### v3

@@ A Multi-Sensory Guidance System for the Visually Impaired Using YOLO and ORB-SLAM
An indoor wayfinding system based on geometric features aided graph SLAM for the visually impaired
Zhang, He and Ye, Cang
IEEE Transactions on Neural Systems and Rehabilitation Engineering
2017
PubMed
[78]
@@
Plane-Aided Visual-Inertial Odometry for 6-DOF Pose Estimation of a Robotic Navigation Aid
Zhang, He and Ye, Cang
IEEE Access
2020
Scopus
[79]
SRAVIP: Smart Robot Assistant for Visually Impaired Persons
Albogamy, Fahad and Alotaibi, Turk and Alhawdan, Ghalib and Mohammed, Faisal
International Journal of Advanced Computer Science and Applications
2021
Forward Snowballing
[80]
@@
A Lightweight Approach to Localization for Blind and Visually Impaired Travelers
Crabb, Ryan and Cheraghi, Seyed Ali and Coughlan, James M
Sensors
2023
Forward Snowballing
[82]
@@ A Lightweight Approach to Localization for Blind and Visually Impaired Travelers
Wearable system to guide crosswalk navigation for people with visual impairment
Son, Hojun and Weiland, James
Frontiers in Electronics
2022
Forward Snowballing
[83]
@@
Indoor Low Cost Assistive Device using 2D SLAM Based on LiDAR for Visually Impaired People
Hakim, Heba and Fadhil, Ali
Iraqi Journal for Electrical & Electronic Engineering
2019
Forward Snowballing
TABLE VIII: List of conference papers included in SLR (part 1)
@@
[84]
Efficient Real-Time Localization in Prior Indoor Maps Using Semantic SLAM
Goswami, R. G. and Amith, P. V. and Hari, J. and Dhaygude, A. and Krishnamurthy, P. and Rizzo, J. and Tzes, A. and Khorrami, F.
9th Inter. Conf. on Automation, Robotics and Applications (ICARA)
2023
IEEE Digital Library
[85]
@@
Detect and Approach: Close-Range Navigation Support for People with Blindness and Low Vision
Hao, Yu and Feng, Junchi and Rizzo, John-Ross and Wang, Yao and Fang, Yi
European Conf. on Computer Vision
2022
Springer Link
[87]
PathFinder: Designing a Map-Less Navigation System for Blind People in Unfamiliar Buildings
@@
Kuribayashi, Masaki and Ishihara, Tatsuya and Sato, Daisuke and Vongkulbhisal, Jayakorn and Ram, Karnik and Kayukawa, Seita and Takagi, Hironobu and Morishima, Shigeo and Asakawa, Chieko
CHI Conf. on Human Factors in Computing Systems
2023
ACM Digital Library
[88]
@@
A Novel Perceptive Robotic Cane with Haptic Navigation for Enabling Vision-Independent Participation in the Social Dynamics of Seat Choice
Agrawal, Shivendra and West, Mary Etta and Hayes, Bradley
IEEE Inter. Conf. on Intelligent Robots and Systems
2022
Scopus
[89]
A Multi-Sensory Blind Guidance System Based on YOLO and ORB-SLAM
Rui, Chufan and Liu, Yichen and Shen, Junru and Li, Zhaobin and Xie, Zaipeng
IEEE Inter. Conf. on Progress in Informatics and Computing (PIC)
2021
IEEE Digital Library
[90]
Indoor Navigation Assistance for Visually Impaired People via Dynamic SLAM and Panoptic Segmentation with an RGB-D Sensor
@@ A Multi-Sensory Blind Guidance System Based on YOLO and ORB-SLAM
Ou, Wenyan and Zhang, Jiaming and Peng, Kunyu and Yang, Kailun and Jaworek, Gerhard and Müller, Karin and Stiefelhagen, Rainer
Inter. Conf. on Computers Helping People with Special Needs
2022
Scopus
[91]
@@
A Wearable Robotic Device for Assistive Navigation and Object Manipulation
Jin, Lingqiu and Zhang, He and Ye, Cang
IEEE Inter. Conf. on Intelligent Robots and Systems
2021
Scopus
[92]
@@ A Wearable Robotic Device for Assistive Navigation and Object Manipulation
Multi-functional smart E-glasses for vision-based indoor navigation
Xu, Jiaqi and Xia, Haisheng and Liu, Yueyue and Li, Zhijun
Inter. Conf. on Advanced Robotics and Mechatronics (ICARM)
2021
Scopus
[93]
Personalized Navigation that Links Speaker’s Ambiguous Descriptions to Indoor Objects for Low Vision People
Lu, Jun-Li and Osone, Hiroyuki and Shitara, Akihisa and Iijima, Ryo and Ryskeldiev, Bektur and Sarcar, Sayan and Ochiai, Yoichi
Inter. Conf. on Human-Computer Interaction
2021
Springer Link
[95]
Guiding Blind Pedestrians in Public Spaces by Understanding Walking Behavior of Nearby Pedestrians
Kayukawa, Seita and Ishihara, Tatsuya and Takagi, Hironobu and Morishima, Shigeo and Asakawa, Chieko
Proc. ACM Interact. Mob. Wearable Ubiquitous Technol.
2020
ACM Digital Library
[96]
A Navigation Aid for Blind People Based on Visual Simultaneous Localization and Mapping
Chen, Cing-Han and Wang, Chien-Chun and Lin, Sian-Fong
IEEE Inter. Conf. on Consumer Electronics
2020
IEEE Digital Library
[97]
Can we unify perception and localization in assisted navigation? an indoor semantic visual positioning system for visually impaired people
Chen, Haoye and Zhang, Yingzhi and Yang, Kailun and Martinez, Manuel and Müller, Karin and Stiefelhagen, Rainer
Computers Helping People with Special Needs: 17th Inter. Conf., ICCHP
2020
Scopus
[98]
@@ A Navigation Aid for Blind People Based on Visual Simultaneous Localization and Mapping
Indoor Localization for Visually Impaired Travelers Using Computer Vision on a Smartphone
Fusco, Giovanni and Coughlan, James M.
17th Inter. web for all Conf.
2020
ACM Digital Library
[99]
@@
Human-Robot Interaction for Assisted Wayfinding of a Robotic Navigation Aid for the Blind
Zhang, He and Ye, Cang
Inter. Conf. on Human System Interaction, HSI
2019
Scopus
[100]
@@
A Multi-Sensor Fusion System for Improving Indoor Mobility of the Visually Impaired
Zhao, Yu and Huang, Ran and Hu, Biao
Chinese Automation Congress (CAC)
2019
IEEE Digital Library
[101]
Navigation Agents for the Visually Impaired: A Sidewalk Simulator and Experiments
@@ A Multi-Sensor Fusion System for Improving Indoor Mobility of the Visually Impaired
Weiss, Martin and Chamorro, Simon and Girgis, Roger and Luck, Margaux and Kahou, Samira E. and Cohen, Joseph P. and Nowrouzezahrai, Derek and Precup, Doina and Golemo, Florian and Pal, Chris
Proceedings of Machine Learning Research
2019
Scopus
[102]
@@
Real-time localization and navigation in an indoor environment using monocular camera for visually impaired
Ramesh, Kruthika and Nagananda, S. N. and Ramasangu, Hariharan and Deshpande, Rohini
5th Inter. Conf. on Industrial Engineering and Applications (ICIEA)
2018
IEEE Digital Library
TABLE IX: List of conference papers included in SLR (part 2)
@@
[103]
Indoor Navigation using Text Extraction
Eden, Jake and Kawchak, Thomas and Narayanan, Vijaykrishnan
IEEE Inter. Workshop on Signal Processing Systems (SiPS)
2018
IEEE Digital Library
[104]
Autonomous Scooter Navigation for People with Mobility Challenges
Mulky, Rajath Swaroop and Koganti, Supradeep and Shahi, Sneha and Liu, Kaikai
IEEE Inter. Conf. on Cognitive Computing (ICCC)
2018
IEEE Digital Library
[105]
Localizing people in crosswalks using visual odometry: Preliminary results
Lalonde, Marc and St-Charles, Pierre-Luc and Loupias, Délia and Chapdelaine, Claude and Foucher, Samuel
Inter. Conf. on Pattern Recognition Applications and Methods (ICPRAM)
2018
Scopus
[106]
@@
Plane-aided visual-inertial odometry for pose estimation of a 3D camera based indoor blind navigation
Zhang, He and Ye, Cang
British Machine Vision Conf. (BMVC)
2017
Scopus
[107]
CCNY Smart Cane
@@
Chen, Qingtian and Khan, Muhammad and Tsangouri, Christina and Yang, Christopher and Li, Bing and Xiao, Jizhong and Zhu, Zhigang
IEEE 7th Annual Inter. Conf. on CYBER Technology in Automation, Control, and Intelligent Systems (CYBER)
2017
IEEE Digital Library
[108]
A Cloud and Vision-Based Navigation System Used for Blind People
Bai, Jinqiang and Liu, Dijun and Su, Guobin and Fu, Zhongliang
Inter. Conf. on artificial intelligence, automation and control technologies
2017
ACM Digital Library
[109]
@@ A Cloud and Vision-Based Navigation System Used for Blind People
Indoor positioning and obstacle detection for visually impaired navigation system based on LSD-SLAM
Endo, Yuki and Sato, Kei and Yamashita, Akihiro and Matsubayashi, Katsushi
Inter. Conf. on Biometrics and Kansei Engineering (ICBAKE)
2017
Scopus
[110]
SeeWay: Vision-Language Assistive Navigation for the Visually Impaired
@@
Yang, Zongming and Yang, Liang and Kong, Liren and Wei, Ailin and Leaman, Jesse and Brooks, Johnell and Li, Bing
IEEE Inter. Conf. on Systems, Man, and Cybernetics (SMC)
2022
Forward Snowballing
[111]
The Methods of Visually Impaired Navigating and Obstacle Avoidance
Shahani, Siddharth and Gupta, Nitin
Inter. Conf. on Applied Intelligence and Sustainable Computing (ICAISC)
2023
Forward Snowballing
[112]

### v4

@@ II-C3 Data extraction
[77] An indoor wayfinding system based on geometric features aided graph SLAM for the visually impaired Zhang, He and Ye, Cang IEEE Transactions on Neural Systems and Rehabilitation Engineering 2017 PubMed
@@
[78] Plane-Aided Visual-Inertial Odometry for 6-DOF Pose Estimation of a Robotic Navigation Aid Zhang, He and Ye, Cang IEEE Access 2020 Scopus
[79] SRAVIP: Smart Robot Assistant for Visually Impaired Persons Albogamy, Fahad and Alotaibi, Turk and Alhawdan, Ghalib and Mohammed, Faisal International Journal of Advanced Computer Science and Applications 2021 Forward Snowballing
@@
[80] A Lightweight Approach to Localization for Blind and Visually Impaired Travelers Crabb, Ryan and Cheraghi, Seyed Ali and Coughlan, James M Sensors 2023 Forward Snowballing
@@
[82] Wearable system to guide crosswalk navigation for people with visual impairment Son, Hojun and Weiland, James Frontiers in Electronics 2022 Forward Snowballing
@@
[83] Indoor Low Cost Assistive Device using 2D SLAM Based on LiDAR for Visually Impaired People Hakim, Heba and Fadhil, Ali Iraqi Journal for Electrical & Electronic Engineering 2019 Forward Snowballing
TABLE X: List of conference papers included in SLR - part 1.
@@
[84] Efficient Real-Time Localization in Prior Indoor Maps Using Semantic SLAM Goswami, R. G. and Amith, P. V. and Hari, J. and Dhaygude, A. and Krishnamurthy, P. and Rizzo, J. and Tzes, A. and Khorrami, F. 9th Inter. Conf. on Automation, Robotics and Applications (ICARA) 2023 IEEE Digital Library
@@
[85] Detect and Approach: Close-Range Navigation Support for People with Blindness and Low Vision Hao, Yu and Feng, Junchi and Rizzo, John-Ross and Wang, Yao and Fang, Yi European Conf. on Computer Vision 2022 Springer Link
@@
[87] PathFinder: Designing a Map-Less Navigation System for Blind People in Unfamiliar Buildings Kuribayashi, Masaki and Ishihara, Tatsuya and Sato, Daisuke and Vongkulbhisal, Jayakorn and Ram, Karnik and Kayukawa, Seita and Takagi, Hironobu and Morishima, Shigeo and Asakawa, Chieko CHI Conf. on Human Factors in Computing Systems 2023 ACM Digital Library
@@
[88] A Novel Perceptive Robotic Cane with Haptic Navigation for Enabling Vision-Independent Participation in the Social Dynamics of Seat Choice Agrawal, Shivendra and West, Mary Etta and Hayes, Bradley IEEE Inter. Conf. on Intelligent Robots and Systems 2022 Scopus
[89] A Multi-Sensory Blind Guidance System Based on YOLO and ORB-SLAM Rui, Chufan and Liu, Yichen and Shen, Junru and Li, Zhaobin and Xie, Zaipeng IEEE Inter. Conf. on Progress in Informatics and Computing (PIC) 2021 IEEE Digital Library
@@
[90] Indoor Navigation Assistance for Visually Impaired People via Dynamic SLAM and Panoptic Segmentation with an RGB-D Sensor Ou, Wenyan and Zhang, Jiaming and Peng, Kunyu and Yang, Kailun and Jaworek, Gerhard and Müller, Karin and Stiefelhagen, Rainer Inter. Conf. on Computers Helping People with Special Needs 2022 Scopus
@@
[91] A Wearable Robotic Device for Assistive Navigation and Object Manipulation Jin, Lingqiu and Zhang, He and Ye, Cang IEEE Inter. Conf. on Intelligent Robots and Systems 2021 Scopus
@@
[92] Multi-functional smart E-glasses for vision-based indoor navigation Xu, Jiaqi and Xia, Haisheng and Liu, Yueyue and Li, Zhijun Inter. Conf. on Advanced Robotics and Mechatronics (ICARM) 2021 Scopus
[93] Personalized Navigation that Links Speaker’s Ambiguous Descriptions to Indoor Objects for Low Vision People Lu, Jun-Li and Osone, Hiroyuki and Shitara, Akihisa and Iijima, Ryo and Ryskeldiev, Bektur and Sarcar, Sayan and Ochiai, Yoichi Inter. Conf. on Human-Computer Interaction 2021 Springer Link
[95] Guiding Blind Pedestrians in Public Spaces by Understanding Walking Behavior of Nearby Pedestrians Kayukawa, Seita and Ishihara, Tatsuya and Takagi, Hironobu and Morishima, Shigeo and Asakawa, Chieko Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 2020 ACM Digital Library
[96] A Navigation Aid for Blind People Based on Visual Simultaneous Localization and Mapping Chen, Cing-Han and Wang, Chien-Chun and Lin, Sian-Fong IEEE Inter. Conf. on Consumer Electronics 2020 IEEE Digital Library
[97] Can we unify perception and localization in assisted navigation? an indoor semantic visual positioning system for visually impaired people Chen, Haoye and Zhang, Yingzhi and Yang, Kailun and Martinez, Manuel and Müller, Karin and Stiefelhagen, Rainer Computers Helping People with Special Needs: 17th Inter. Conf., ICCHP 2020 Scopus
@@
[98] Indoor Localization for Visually Impaired Travelers Using Computer Vision on a Smartphone Fusco, Giovanni and Coughlan, James M. 17th Inter. web for all Conf. 2020 ACM Digital Library
@@
[99] Human-Robot Interaction for Assisted Wayfinding of a Robotic Navigation Aid for the Blind Zhang, He and Ye, Cang Inter. Conf. on Human System Interaction, HSI 2019 Scopus
@@
[100] A Multi-Sensor Fusion System for Improving Indoor Mobility of the Visually Impaired Zhao, Yu and Huang, Ran and Hu, Biao Chinese Automation Congress (CAC) 2019 IEEE Digital Library
@@
[101] Navigation Agents for the Visually Impaired: A Sidewalk Simulator and Experiments Weiss, Martin and Chamorro, Simon and Girgis, Roger and Luck, Margaux and Kahou, Samira E. and Cohen, Joseph P. and Nowrouzezahrai, Derek and Precup, Doina and Golemo, Florian and Pal, Chris Proceedings of Machine Learning Research 2019 Scopus
@@
[102] Real-time localization and navigation in an indoor environment using monocular camera for visually impaired Ramesh, Kruthika and Nagananda, S. N. and Ramasangu, Hariharan and Deshpande, Rohini 5th Inter. Conf. on Industrial Engineering and Applications (ICIEA) 2018 IEEE Digital Library
TABLE XI: List of conference papers included in SLR - part 2.
@@
[103] Indoor Navigation using Text Extraction Eden, Jake and Kawchak, Thomas and Narayanan, Vijaykrishnan IEEE Inter. Workshop on Signal Processing Systems (SiPS) 2018 IEEE Digital Library
[104] Autonomous Scooter Navigation for People with Mobility Challenges Mulky, Rajath Swaroop and Koganti, Supradeep and Shahi, Sneha and Liu, Kaikai IEEE Inter. Conf. on Cognitive Computing (ICCC) 2018 IEEE Digital Library
[105] Localizing people in crosswalks using visual odometry: Preliminary results Lalonde, Marc and St-Charles, Pierre-Luc and Loupias, Délia and Chapdelaine, Claude and Foucher, Samuel Inter. Conf. on Pattern Recognition Applications and Methods (ICPRAM) 2018 Scopus
@@
[106] Plane-aided visual-inertial odometry for pose estimation of a 3D camera based indoor blind navigation Zhang, He and Ye, Cang British Machine Vision Conf. (BMVC) 2017 Scopus
@@
[107] CCNY Smart Cane Chen, Qingtian and Khan, Muhammad and Tsangouri, Christina and Yang, Christopher and Li, Bing and Xiao, Jizhong and Zhu, Zhigang IEEE 7th Annual Inter. Conf. on CYBER Technology in Automation, Control, and Intelligent Systems (CYBER) 2017 IEEE Digital Library
[108] A Cloud and Vision-Based Navigation System Used for Blind People Bai, Jinqiang and Liu, Dijun and Su, Guobin and Fu, Zhongliang Inter. Conf. on artificial intelligence, automation and control technologies 2017 ACM Digital Library
@@
[109] Indoor positioning and obstacle detection for visually impaired navigation system based on LSD-SLAM Endo, Yuki and Sato, Kei and Yamashita, Akihiro and Matsubayashi, Katsushi Inter. Conf. on Biometrics and Kansei Engineering (ICBAKE) 2017 Scopus
@@
[110] SeeWay: Vision-Language Assistive Navigation for the Visually Impaired Yang, Zongming and Yang, Liang and Kong, Liren and Wei, Ailin and Leaman, Jesse and Brooks, Johnell and Li, Bing IEEE Inter. Conf. on Systems, Man, and Cybernetics (SMC) 2022 Forward Snowballing
[111] The Methods of Visually Impaired Navigating and Obstacle Avoidance Shahani, Siddharth and Gupta, Nitin Inter. Conf. on Applied Intelligence and Sustainable Computing (ICAISC) 2023 Forward Snowballing